# Edge of Stochastic Stability: Revisiting the Edge of Stability for SGD

### v3

@@ 5 The Geometry of the Landscape at EoSS
Then, we establish here that:
  • Also $\lambda_{\max}^{b}$ stabilizes.
  • $\lambda_{\max}^{b}$ is always greater than $\lambda_{\max}$.
  • However, the gap between $\lambda_{\max}^{b}$ and $\lambda_{\max}$ typically goes as $1/b^{\alpha}$ generally with $\alpha=0.70$. This is surprising, as when fixing the network, the gap goes as $1/b$ or $1/b^{1/2}$ both empirically and mathematically.
5.1 Average $\lambda_{\max}^{b}$ for the mini-batch Hessians
We establish that $\lambda_{\max}^{b}$–the average over the possible batches sampled of the highest eigenvalue of the mini-batch Hessian–generally stabilizes at a value just bigger than $2/\eta$. This happens on a wide range or models and datasets.
Interestingly, we observe that $\lambda_{\max}^{b}$ responds to the change of parameters exactly as Cohen et al. (2021) observed for $\lambda_{\max}$ in the full-batch setting.
5.2 Size of $\lambda_{\max}$
Although we show that the quantity of interest is the mini-batch Hessian, in case of mini-batch SGD, there are both historical and generalization-related reasons to investigate what the spectrum of the full-batch Hessian is.
We establish that also $\lambda_{\max}$ generally stabilizes in the EoSS at a level which is strictly smaller than $2/\eta$. We show that the exact level at which $\lambda_{\max}$ does not only depend on learning rate and batch size but on the full trajectory taken, unlike for EoS (Cohen et al., 2021). Indeed, in Section 3.3 we showed that it does not fully adapt to changes in hyper parameters. However, we observed that when we do not induce any shock in the system by changing hyper parameters during training, we can characterize two regimes for the stabilization levels, see Figure 6. Precisely
  • (i) $b\leq$ critical batch size: $\lambda_{\max}$ grows with $b$ with a power law.
  • (ii) $b\geq$ critical batch size: $\lambda_{\max}$ grows with $b$ much slower while approaching $2/\eta$ from below.
Unfortunately, we could not find a way to relate the size of $\lambda_{\max}$ to learning rate and batch size–see Figure 9–if not through the gap it has with the average $\lambda_{\max}$ of the mini-batch Hessian $\lambda_{\max}^{b}$. Notably, the critical batch size we mention is the point beyond which training enters the full-batch regime. Works as Zhang et al. (2024) study the following notion of critical batch size: "the point beyond which increasing batch size may result in computational efficiency degradation". We conjecture there may be a relation between the two quantities and we leave it to future work.
5.3 The Gap Between $\lambda_{\max}$ and $\lambda_{\max}^{b}$
Note that conceptually $\lambda_{\max}^{b}$ is inherently of the same nature as $\lambda_{\max}$, with the only difference being the "part" of the loss it is computed on-the mini-batch loss for the former, and full-batch loss for the latter. Yet, this difference proves to be crucial in being descriptive of the SGD training dynamics. Indeed, e.g.: {mdframed}
While the mini-batch gradients are an unbiased estimator of the full-batch gradient, $\lambda_{\max}\big{(}\mathcal{H}(L_{B})\big{)}$ is a positively biased estimator of $\lambda_{\max}$.
We indeed show in Lemma 4, that $\lambda_{\max}^{b}$ is strictly bigger than $\lambda_{\max}$ and, as a consequence, decreasing the batch size the gap increases. In particular,
Figure 6: Batch Sharpness vs. $\lambda_{\max}^{b}$ vs. $\lambda_{\max}$ at EoSS. The values of Batch Sharpness (green), mini-batch $\lambda_{\max}^{b}$ (orange), and full-batch $\lambda_{\max}$ (blue) against varying batch sizes.
We now empirically show what the gaps $\lambda_{\max}^{b}-\lambda_{\max}$ are, and we discuss why that is surprising. As shown in Appendix G and Figure 9, the log-log plot of the gap against batch size reveals three distinct regimes:
  • Very small $b$: for extremely small batch sizes, the gap plateaus. This is unsurprising, as we are in the (i) regime for $\lambda_{\max}$.
  • Typical/moderate $b$: the gap consistently scales as $1/b^{0.7}$ in the experiment. This trend dominates for a wide range of practical batch sizes, highlighting the systematic scaling behavior of the $\lambda_{\max}^{b}-\lambda_{\max}$ gap. It is important to note that we establish in Appendix G the exponent to be $0.7$ independently from the depth, the seeds, the subsets of Cifar10 considered, and the architecture (MLP, CNN, ResNets). Also note that the critical batch size falls within this range. The limitation of this observation in describing $\lambda_{\max}$ is that this consistent trend is consequence both of $\lambda_{\max}$ growing and $\lambda_{\max}^{b}$ decreasing, not just one of the two.
  • Very big $b$: here generally $\lambda_{\max}^{b}$ is essentially $\lambda_{\max}$ and the gap is converging to zero, we conjecture that some form of scaling limit from random matrix theory applies here. The scaling law is generally of the form: $1/b^{\alpha}$ with $\alpha\geq 1$. The exact value of $\alpha$ may vary with the setting, but the overall trend is robust.
While the behavior for very small batches is not surprising, for bigger batch sizes it is. In computer vision tasks where there are way more parameters than datapoints, we observe that the gap between $\lambda_{\max}^{b}$ and $\lambda_{\max}$ decreases linearly with $1/b$ when evaluating the $\lambda_{\max}^{b}$ of any fixed model with different batch sizes, see Figure 8. This $1/b$ scaling is also what one expects from the analysis in Appendix F.
6 Implications: Why Noise-Injected GD Differs from Mini-Batch SGD
6.1 SGD vs. Noisy Gradient Descent
  • Directional Noise: Gaussian noise matching the directions and signs of mini-batch SGD noise (Simsekli et al., 2019).
  • Diagonal Noise: Gaussian noise restricted to the diagonal part of the SGD noise covariance.
  • Isotropic Noise: Gaussian noise with isotropic covariance.
As shown in Fig. 7, only directionally accurate noise leads to an EoSS-like regime with $\lambda_{\max}$ stabilizing well below $2/\eta$. More generic (diagonal or isotropic) noise fails to reproduce this behavior, indicating that standard SDE approximations—which typically ignore directional correlations—cannot fully capture the observed dynamics.
Figure 7: SGD vs. Noisy GD. Only noise preserving the covariance structure and direction of SGD leads to $\lambda_{\max}$ plateauing below $2/\eta$ (akin to EoSS). Diagonal or isotropic noise fails to reproduce this behavior. Note that batch sharpness is still plotted for comparison, though for noisy GD it loses its meaning.
From a theoretical perspective (Appendix D), these experiments suggest that stability thresholds differ fundamentally between mini-batch SGD (governed by Batch Sharpness) and noise-injected GD (governed by $\lambda_{\max}$).
6.2 The Mini-Batch Landscape Matters
Classical analyses of neural network optimization often assume a single, static landscape: (i) Online perspective, modeling each sample’s gradient as a noisy unbiased estimator of the expected risk, or (ii) Offline perspective, treating the dataset as fixed and SGD as noisy GD on the empirical loss. In both views, it is the full-batch Hessian that supposedly drives curvature. Our results instead highlight the mini-batch landscape: each update sees a Hessian $\mathcal{H}(L_{B})$ that can differ significantly from $\mathcal{H}(L)$, leading to Batch Sharpness stabilizing at $2/\eta$ even when full-batch $\lambda_{\max}$ is much smaller.
6.3 Challenges for SDE Modeling
@@
Prior works already note limitations of SDE-based approaches for SGD implicit regularization: they may be mathematically ill-posed (Yaida, 2018), fail except under restrictive conditions (Li et al., 2021), converge to qualitatively different minima (HaoChen et al., 2020), or miss higher-order effects (Damian et al., 2021; Li et al., 2022). Recent discrete analyses (Roberts, 2021; Smith et al., 2021; Beneventano, 2023) attempt to address some of these issues. Nonetheless, our findings expose a deeper gap: when batch sizes are small, the geometry of the mini-batch Hessian differs markedly from that of the full-batch, altering both eigenvalues and eigenvector alignments. Conventional SDE models, which assume a static or average Hessian, cannot easily capture these rapid fluctuations. Fully accounting for the distribution of $\{\mathbf{A}_{i}\}$ remains an open challenge, emphasizing the need for further research into truly batch-specific dynamics.
@@ Conclusions.
We introduced Batch Sharpness as a key descriptor of mini-batch SGD dynamics, showing that it stabilizes near $2/\eta$. This defines an Edge of Stochastic Stability, analogous to the classic EoS (Xing et al., 2018; Cohen et al., 2021, 2022), but now in the small-batch regime most commonly used in practice. Empirically, Batch Sharpness remains at $2/\eta$ while the full-batch $\lambda_{\max}$ hovers below, at a level that systematically depends on batch size; indeed, the gap $\lambda_{\max}^{b}-\lambda_{\max}$ follows a power law $(\text{batch size})^{-0.7}$ across various models and datasets.
@@
(i) We have tested only image-classification tasks, leaving open whether similar phenomena arise in NLP, RL, or other domains. (ii) Although we observe empirical batch-size scaling laws, we lack a full theoretical account, especially at moderate batch sizes. (iii) Our experiments mainly use fixed learning rates and standard architectures, so very large-scale or large-batch settings remain less explored. (iv) We have not analyzed momentum-based or adaptive methods (e.g. Adam), even though full-batch EoS has been seen there (Cohen et al., 2022).
@@
Beyond addressing these limitations, several directions remain: (i) Extending higher-order analyses (Damian et al., 2023; Cohen et al., 2021; Ahn et al., 2022; Lyu et al., 2023) to mini-batch SGD may clarify why $\textit{Batch Sharpness}\sim 2/\eta$, (ii) Refining our batch-size scaling argument via random matrix theory might explain the “typical” exponent (around 0.7) and its possible crossovers for very large or very small batches.
In summary, Batch Sharpness provides a more faithful perspective on SGD than the full-batch $\lambda_{\max}$. Our findings highlight how mini-batch Hessians drive training instabilities and ultimately shape learned solutions, hopefully inspiring further research on how step size and batch size jointly govern deep network training.
@@ Acknowledgement
A special thanks to Stanisław Jastrzębski, Alex Damian, Jeremy Cohen, Afonso Bandeira, Boris Hanin, Mark Lowell, and Renée Carmona for invaluable discussions, which were crucial to the development of this project. We acknowledge the use of DeepSeek, Claude, and ChatGPT for providing code assistance, debugging support, and editing suggestions.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A On the Stability
A.1 On Equivalent Formulations of Stability
Consider as working example the quadratic Taylor approximation of a loss function:
$f(x)=\frac{1}{2}x^{\top}\mathbf{A}x+\mathbf{b}^{\top}x+\mathbf{c},$
where $\mathbf{A}$ is the Hessian matrix (symmetric and positive semidefinite), $\mathbf{b}$ is the gradient vector, and $\mathbf{c}$ is a constant.
The question of stability is very nuanced, and we decided to define stability and $\eta_{\max}$ as follows
Definition 2 (Stability).
We say that the optimization algorithm $\mathcal{A}$ on the objective $f$ is stable if it does not diverge.
By Folklore, in the quadratic objective $f$ above, enforcing stability is equivalent to choosing a step size smaller than $\eta_{\max}$ as follows.
Definition 3 (Step size at the edge for an algorithm).
Specify a discrete-time optimization algorithm of choice $\mathcal{A}$. We define "step size at the edge for $\mathcal{A}$" the maximum step size $\eta_{\max}(\mathcal{A})>0$ under which the dynamics do not diverge along the whole trajectory in the quadratic objective case, meaning there does not exist a $t$ such that $|x_{t+1}-x^{*}|>|x_{t}-x^{*}|$. In case of a stochastic update, we define $\eta_{\max}(\mathcal{A})$ analogously, but in expectation.
This is a good definition in the quadratic setting above, no matter how we define divergence, indeed
Lemma 1.
Let $f(x)$ be quadratic, specify a discrete-time optimization algorithm of choice $\mathcal{A}$. Then $\eta_{\max}(\mathcal{A})>0$ is the maximum step size under which the dynamics does not diverge to infinity. Both in terms of $f(x_{t})$ and in terms of $\|x_{t}-x^{*}\|$.
And we have, analogously, stability in terms of non-increasing loss.
Lemma 2.
Let $f(x)$ be quadratic, specify a discrete-time optimization algorithm of choice $\mathcal{A}$. Then $\eta_{\max}(\mathcal{A})>0$ is the maximum step size under which there does not exist a $t$ such that $f(x_{t+1})>f(x_{t})$.
A.2 Specialization of the Setting to Mini-Batch SGD
In the case of mini-batch SGD, we are working, instead, with the following 2-order Taylor approximation:
$f_{B}(x)=\frac{1}{b}\sum_{i\in B}\left(\frac{1}{2}x^{\top}\mathbf{A}_{i}x+\mathbf{b}_{i}^{\top}x+\mathbf{c}_{i}\right),$ (6)
where $B\subset\{1,2,\ldots,n\}$ is a randomly sampled mini batch of size $b$, and $\mathbf{A}_{i}$, $\mathbf{b}_{i}$, and $\mathbf{c}_{i}$ represent contributions from individual data points. The full objective matrices are instead given by:
$\mathbf{A}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{A}_{i},\quad\mathbf{b}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{b}_{i}.$
Update rule for mini-batch SGD.
Here, the mini-batch SGD update rule is:
$x_{t+1}=x_{t}-\eta\nabla f_{B}(x_{t}),$
where the gradient for mini batch $B$ is:
$\nabla f_{B}(x_{t})\ =\ \frac{1}{b}\sum_{i\in B}\left(\mathbf{A}_{i}x_{t}+\mathbf{b}_{i}\right)\ =\ \left[\frac{1}{b}\sum_{i\in B}\mathbf{A}_{i}\right](x_{t}-x^{*})\,+\,\left(\frac{1}{b}\sum_{i\in B}\mathbf{A}_{i}x^{*}+\mathbf{b}_{i}\right).$ (7)
Substituting this gradient, the update becomes:
$x_{t+1}=x_{t}-\eta\left(\frac{1}{b}\sum_{i\in B}\mathbf{A}_{i}x_{t}+\frac{1}{b}\sum_{i\in B}\mathbf{b}_{i}\right).$
To analyze stability, one generally focuses on the evolution of the error $x_{t}-x^{*}$, where $x^{*}=-\mathbf{A}^{-1}\mathbf{b}$ is the optimum. By adding and subtracting $x^{*}$, we obtain that error evolves as:
$x_{t+1}-x^{*}=\left(I-\eta\frac{1}{b}\sum_{i\in B}\mathbf{A}_{i}\right)\big{(}x_{t}-x^{*}\big{)}\ -\ \eta\left(\sum_{i\in B}\mathbf{A}_{i}x^{*}+\frac{1}{b}\sum_{i\in B}\mathbf{b}_{i}\right).$
We can notice here the first difference with the full-batch case:
@@ A.2 Specialization of the Setting to Mini-Batch SGD
    The rest $\sum_{i\in B}\mathbf{A}_{i}x^{*}+\frac{1}{b}\sum_{i\in B}\mathbf{b}_{i}$ it is different than 0 and it is different than 0 even in expectation when sampling is not uniform and independent.
@@
    The multiplier $\frac{1}{b}\sum_{i\in B}\mathbf{A}_{i}$ is not $\mathbf{A}$.
Next note that if we work with an full-batch GD with noise injected with unbiased and independent steps in the fashion of Robbins & Monro (1951), their expectation evens these terms out. We refer the reader to Appendix D for that. If we work with a biased algorithm, as Random Reshuffling (or SGD without replacement) or Shuffle Once, this is not the case (Mishchenko et al., 2020; Smith et al., 2021; Beneventano, 2023). We refer the reader to Appendix E for what changes in this setting.
Furthermore, while in the deterministic case working with $x_{t+1}-x^{*}$ or with $\|x_{t+1}-x^{*}\|^{2}$ is the same, this is not the case for mini-batch algorithms. It is thus non-obvious how to generalize stability to start with. We will see in the next section how to do it.
A.3 Formulating Stability for Mini-Batch SGD
Unlike in the deterministic case, we have to be a little more careful with the formulation here. Let $(\lambda_{i},v_{i})$ be the ordered eigenvalue-eigenvector pairs of $\mathbf{A}$. With a deterministic Hessian matrix $\mathbf{A}$ we have that $x_{t+1}-x^{*}$ does not diverge if and only if the highest eigenvalue $\alpha_{i}$ such that $\langle x_{t+1}-x^{*},v_{i}\rangle\neq 0$ is smaller than $2/\eta$, see Cohen et al. (2021). The idea is that if a component $\langle x_{t+1}-x^{*},v_{i}\rangle\neq 0$ exists, after $k$ steps it gets multiplied by $(1-\eta\alpha_{i})^{k}$. This implies that the system will diverge if and only if $|1-\eta\alpha_{i}|>1$. When our step is a noise-injected GD similar stability phenomena happens, $\mathbf{A}$ is not impacted by the noise, see Appendix D.
However, in the case of mini-batch SGD we have to be much more precise, and note that at every step, stability does not depend on $\mathbf{A}$, but on the "$\mathbf{A}$" of the mini batch. We have an important changes here:
We do not have persistence of a Hessian matrix "$\mathbf{A}$". In particular, there is no persistence in the eigenspaces of the Hessian from step to step. This implies that we can not make an argument as the one above tracking the components of $x_{t}-x^{*}$. This will have an important implication, that at the step $t$ we care about the direction we are moving towards, not about the eigenvector of the highest eigenvalue. We are now dealing with objects of the form $\nabla f^{\top}\cdot\nabla^{2}f\cdot\nabla f$, not of the form $\lambda_{\max}\nabla^{2}f$.
This imply that we may not really work on the eigenspaces of $\mathbf{A}$ as before, we decide to work with the analogous formulation of stability given by Lemma 2: For quadratic objectives, $|x_{t+1}-x^{*}|\leq|x_{t}-x^{*}|$ reduces in norm if and only if $f(x_{t+1})\leq f(x_{t})$. We will work from now on with the formulation of stability of $f(x_{t+1})\leq f(x_{t})$.
At the step $t$ we can write the following thing update for a quadratic objective $f$.

### v4

@@ 7 Implications: How Noise-Injected GD Differs from SGD
Standard SDE—or analogous—approximations of SGD cannot thus describe the location of convergence of SGD or its behavior for neural networks under the assumption of progressive sharpening. Indeed, they typically ignore any statistics of the Hessians except for the mean.
@@
Prior works already note limitations of SDE-based approaches for SGD implicit regularization: they may be mathematically ill-posed [53], fail except under restrictive conditions [54], converge to qualitatively different minima [55], or miss higher-order effects [56, 57]. Recent discrete analyses [36, 37, 58] attempt to address some of these issues. Nonetheless, our findings expose a deeper gap: when batch sizes are small, the geometry of the mini-batch Hessian differs markedly from that of the full-batch, altering both eigenvalues and eigenvector alignments. Conventional SDE models, which assume a static or average Hessian, cannot easily capture these rapid fluctuations.
@@ Conclusions.
We have addressed the longstanding question of if and how mini-batch SGD enters a regime reminiscent of the “Edge of Stability” previously observed in full-batch methods. Contrary to the usual focus on the global Hessian’s top eigenvalue, we uncovered that Batch Sharpness—the expected directional curvature of the mini-batch Hessian in the direction of its own gradient—consistently rises (progressive sharpening) and then hovers around $2/\eta$, independent of batch size. This behavior characterizes a new “Edge of Stochastic Stability”, which explains how mini-batch training can exhibit catapult-like surges and settle into flatter minima even when the full-batch Hessian remains below $2/\eta$. Our analysis clarifies why smaller batch sizes and larger learning rates both constrain the final curvature to a lower level, thereby linking these hyperparameters to flatter solutions and often improved generalization. Furthermore, we show that this phenomenon depends on the directional structure of mini-batch noise and does not straightforwardly emerge from generic isotropic or diagonal noise injections, highlighting important limitations of SDE-based approximations. Overall, the EoSS framework unifies several empirically observed effects—catapult phases, dependence on batch size, and progressive sharpening—under a single perspective focused on the mini-batch landscape and its directional curvature.
@@
(i) We have tested only image-classification tasks, leaving open whether similar phenomena arise in NLP, RL, or other domains. (ii) Although we observe empirical EoSS, we lack full theoretical account. (iii) Our experiments mainly use fixed learning rates and standard architectures, so very large-scale or large-batch settings remain less explored. (iv) We have not analyzed momentum-based or adaptive methods (e.g. Adam), even though full-batch EoS has been seen there [3].
@@
Beyond addressing these limitations, several directions remain: Understanding (i) where $\lambda_{\max}$ stabilizes; (ii) how EoSS and EoS affect performances and the features learned by the neural network [1, 18, 21, 24]; (iii) consequently if it is benign effect or not; (iv) what the other sources of instability are there in the training; (v) better describing the phenomenon of progressive sharpening; and (vi) understanding its causes. Moreover, we conjecture, and experiments seem to suggest, that the phenomenon for SGD is very similar to the one described by [18, 26], but a mathematical verification is a necessary future work.
@@ Acknowledgement
A special thanks to Stanisław Jastrzębski, Alex Damian, Jeremy Cohen, Afonso Bandeira, Boris Hanin, and Renée Carmona for invaluable discussions, which were crucial to the development of this project. We also want to thank Mark Lowell, Yee Whye Teh, and the referees of ICML 2025, for comments and feedbacks on the first version of the preprint. We acknowledge the use of DeepSeek, Claude, and ChatGPT for providing code assistance, debugging support, and editing suggestions.
Appendix A On the Two Types of Oscillations in SGD Dynamics
A fundamental challenge in analyzing SGD compared to GD stems from the inherent oscillations induced by mini-batch gradient noise. This appendix, together with Appendix C, extends the discussion in Section 3 by formally distinguishing between two distinct types of oscillations: noise-driven (Type-1) and curvature-driven (Type-2). This distinction is crucial because Type-1 oscillations occur independently of the loss landscape’s curvature and thus do not exert a regularizing effect on the sharpness of the final solution. In contrast, Type-2 oscillations are directly caused by landscape curvature and induce an implicit regularization effect by discouraging convergence towards sharp minima.
We begin with a minimalistic example to illustrate the nature of Type-1
A.1 A Minimalistic Quadratic Example.
Consider a regression problem with two datapoints, $1$ and $-1$, and a linear model $f(x)=x$ under the quadratic loss. The (scaled) full-batch loss is given by:
$L(x)\;=\;\frac{1}{4}(x-1)^{2}\;+\;\frac{1}{4}(x+1)^{2}.$
Batch-1 SGD updates with step-size $20<\eta<$ 2 result in oscillatory behavior around the optimum $x=0$ due entirely to gradient noise, with amplitude approximately $\sqrt{\frac{\eta}{2-\eta}}$. Crucially, the Hessian in this example is small ($\frac{d^{2}L}{dx^{2}}=1$), demonstrating that these persistent oscillations are entirely noise-driven (Type-1).
Formally, the SGD update is:
$x_{t+1}=x_{t}-\eta\nabla\ell_{i_{t}}=(1-\eta)x_{t}+\eta\xi_{t}$
where $l_{i_{t}}$s are the individual datapoint losses, and $\xi_{t}$s are i.i.d Rademacher random variables. Thus, we obtain the first two moments explicitly:
$\mathbb{E}[x_{t}]=(1-\eta)\mathbb{E}[x_{t-1}]=(1-\eta)^{t}x_{0}$
$\mathbb{E}[x_{t}^{2}]=(1-\eta)^{2}\mathbb{E}[x_{t-1}^{2}]+\eta^{2}=(1-\eta)^{2t}x_{0}+\frac{\eta^{2}}{1-(1-\eta)^{2}}\Bigl{(}1-(1-\eta)^{2t}\Bigr{)}$
This implies convergence in expectation for $20<\eta<$ 2, with a limiting variance given by:
$\lim_{t\to\infty}\mathbb{E}[x_{t}^{2}]=\frac{\eta}{2-\eta}$
and divergence for $\eta>2$.
A key observation is that increasing $\eta$ to any value $\eta_{1}<2$ merely changes the amplitude of oscillations to $\sqrt{\frac{\eta_{1}}{2-\eta_{1}}}$ without triggering any catapult-like behavior. The only step size for which we observe Type-2 (curvature-driven) oscillations and an EoS-like [^8] The key difference between these oscillations and genuine EoS behavior in neural networks is that, in the quadratic case, the full-batch loss does not decrease, making this scenario inherently less informative. In contrast, neural networks exhibit a surprising, albeit non-monotonic, decrease in loss within this instability regime, an effect arising from the multidimensional nature of their optimization landscape [18] instability is precisely $\eta=2$, where the dynamics effectively become a random walk, and any larger step size leads to divergence.
See Figure 3 for a plot of this phenomena for a one-dimensional example with many datapoints, with the same stability thresholds (also see Example 3.1 in Bottou et al. [48]).
{mdframed}
Crucially, when $\eta<2$ oscillations occur persistently on the full-batch loss, despite the individual steps on the mini-batch loss remaining stable.
The oscillation is due to the fact that the mini-batch loss landscape shifts from step to step, not to the fact that the steps are unstable.
A.2 General Case: From One-Dimensional Toy to Multidimensional Lyapunov Analysis
The simple one-dimensional regression in §A.1 already demonstrates how noise-driven (Type-1) oscillations can persist indefinitely and yield a stable “two-cycle” around the optimum, independent of the actual Hessian magnitude. In higher dimensions, the story is similar: when the step size $\eta$ is fixed, the randomness in mini-batch gradients still injects a continual “kick” at each iteration, causing the iterates to hover in a noisy neighborhood of the minimum. The main difference is that now there can be many directions—some with higher curvature than others, or even flat ($\lambda=0$) directions. Nonetheless, the essential mechanism remains:
$\Delta_{t+1}\;=\;\bigl{(}I-\eta\,\mathcal{H}\bigr{)}\;\Delta_{t}\;-\;\eta\,\xi_{t},$
where $\Delta_{t}=x_{t}-x^{\star}$ is the displacement from the optimum, $\mathcal{H}$ is the Hessian at $x^{\star}$, and $\xi_{t}$ encodes the random fluctuation (gradient noise). Once $\Delta_{t}$ settles into a stationary distribution, the covariance $\Sigma_{x}$ can be found by solving a discrete Lyapunov equation similar to the one-dimensional case.
Key References and the Road Ahead.
A number of works formalize this “SGD noise equilibrium” by viewing the updates as a linear Markov chain in a neighborhood of $x^{\star}$. Classical references include Mandt et al. [47] for the stochastic differential analogy (Ornstein–Uhlenbeck process), and Bottou et al. [48] for a thorough discussion of how the constant stepsize prevents exact convergence. Intuitively, the argument for Proposition 1 goes thus as follows:
@@ A.2 General Case: From One-Dimensional Toy to Multidimensional Lyapunov Analysis
    The iterates oscillate with a stationary covariance $\Sigma_{x}$ around $x^{*}$.
@@
    Full-batch (expected) gradient is zero at $x^{*}$ and grows roughly linearly with distance for small deviations (by Taylor expansion $\nabla L(x)\approx\mathcal{H}(x-x^{*})$). So on average over the iterations we have
    $\mathbb{E}_{k}[\lVert\nabla L(x_{k})\rVert^{2}]=\mathrm{Tr}\big{(}\mathcal{H}\Sigma_{x}\mathcal{H}).$
  3. 3.
    The stationary covariance of the gradients is $\Sigma_{g}$ satisfies:
    $\Sigma_{x}\approx\frac{\eta}{2}(\mathcal{H}^{-1}\Sigma_{g}).$
Putting all together, this implies that $\Sigma_{g}$ of the gradients is about $\frac{2}{\eta}\mathcal{H}^{-1}$ bigger than the full-batch $\nabla L\nabla L^{\top}$. Precisely the following quantity (where $L_{i}$ is the loss on the $i-th$ data point):
$\frac{\mathbb{E}_{i}\big{[}\nabla L_{i}(\mathbf{x})^{\top}\mathcal{H}\nabla L_{i}(\mathbf{x})\big{]}}{\lVert\nabla L\rVert^{2}}\,=\,\frac{\mathrm{Tr}\big{(}\mathcal{H}\Sigma_{g}\big{)}}{\mathrm{Tr}\big{(}\mathcal{H}\Sigma_{x}\mathcal{H})},$ (3)
and this can be rewritten as